OpenStudy (anonymous):

Can somebody show me how to format this equation in standard form? Input in standard form the equation of the given line.The line with m = -3/2 and b = 6 Can somebody show me how to format this equation in standard form? Input in standard form the equation of the given line.The line with m = -3/2 and b = 6 @Mathematics

OpenStudy (anonymous):

Standard form is Ax+By=C

OpenStudy (anonymous):

So your equation is y = -3/2 x +6 standard form does not have fractions so multiply your equation by 2 2y = -3x + 6 3x+2y=6

OpenStudy (anonymous):

Alright! Thanks... let me see if I can do some on my own

OpenStudy (anonymous):

It says it is wrong!

OpenStudy (anonymous):

oh sorry I forgot to multiply the 6 by 2 3x+2y=12

OpenStudy (anonymous):

Yup...

OpenStudy (anonymous):

Input in standard form the equation of the given line.The line with m = -1 /2 and b = 1

OpenStudy (anonymous):

y-y1=x-x1

OpenStudy (anonymous):

y+4=-3/8x-5

OpenStudy (anonymous):

y+4=-3/8x+-15/8

OpenStudy (anonymous):

equation is y = -1/2 x + 1 2(y = -1/2 x + 1) 2y=-x+2 x + 2y = 2

OpenStudy (anonymous):

So I am wrong with the steps?

OpenStudy (anonymous):

y+4=-3/8x-5 y=-3/8x -9 8(y=-3/8x -9) 8y = -3x -72 3x + 8y = -72

OpenStudy (anonymous):

Ok I was close???

OpenStudy (anonymous):

I'm not really sure where you for -15/8

OpenStudy (anonymous):

It is not accepting it.

OpenStudy (anonymous):

-3/8x-5

OpenStudy (anonymous):

is it  or